OpenStudy (anonymous):

Find the 4th term of the series- 2, 5/2, 10/3,.. ans.5 (How?)

OpenStudy (lgbasallote):

2 + 1/2 = 5/2 5/2 + 2/2 = 7/2 so most probably the next one will be 7/2 + 3/2 = 10/2 then 10/2 + 4/2 = 14/2 = 7

OpenStudy (anonymous):

sorry pls check the ques again.

OpenStudy (lgbasallote):

do you want a formal formula?

OpenStudy (anonymous):

that was 10/3

OpenStudy (lgbasallote):

oh

hartnn (hartnn):

then maybe 17/4...

OpenStudy (lgbasallote):

\[\frac 21 , \; \frac 52, \ \frac{10}{3}\] the third term would probably be \[\frac{10 \times 5}{4} \implies \frac{50}{4}\] then the 4th term would probably be\[\frac{50 \times 10}{5} \implies \frac{500}{5} \implies 100\]

OpenStudy (lgbasallote):

does that make sense?

OpenStudy (anonymous):

the ans. is 5

OpenStudy (lgbasallote):

bummer

hartnn (hartnn):

yes, it comes out to be 5....17/4....then 25/5=5

hartnn (hartnn):

the diffeerence in numerator is 3,5,7,9...

OpenStudy (lgbasallote):

\[\frac 21 , \; \frac 52, \; \frac{10}{3}\] so the 3rd term would be \[\frac{10+7}{4} \implies \frac{17}{4}\] then the 4th term would be \[\frac {17 + 9}{5} \implies \frac{26}{5}\] not 25/5

OpenStudy (anonymous):

note : this ques. is based on sequences (i.e. AP,GP or HP)

OpenStudy (lgbasallote):

lol i got it

OpenStudy (lgbasallote):

\[\frac 21, \; \frac 52, \; \frac{10}{3}\] 4th term would be \[\frac{10 \times 2}{4} \implies \frac{20}{4} \implies 5\]

OpenStudy (lgbasallote):

i forgot 3 terms were already given

OpenStudy (lgbasallote):

although that does not explain 5 :/

OpenStudy (lgbasallote):

i mean 5/3

OpenStudy (anonymous):

OH got it,